OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. Please show all of your work. Write your final answer exactly as a fraction in lowest terms 22/45/ 14/25

OpenStudy (anonymous):

is the 22/45 over the 14/25?

OpenStudy (anonymous):

yes buddy

OpenStudy (anonymous):

22/45 divided by 14/25

OpenStudy (anonymous):

(22/45) / (14/25) = (22/45) x (25/14) cross multiply and simplify.

OpenStudy (anonymous):

sorry, not cross multiply. just multiply them out as fractions and simplify.

OpenStudy (anonymous):

yes they aet over top ts beside jus cant make the diivison sign

OpenStudy (anonymous):

arent*

OpenStudy (anonymous):

you'll come up with (22 x 25) / (45 x 14) = 550 / 350, which simplifies to 11/7

OpenStudy (anonymous):

actually hang on- i think i screwed up somewhere.

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

yeah- NO idea where i came up with that. what you ACTUALLY have to do is just multiply them as fractions once you flip the division into a multiplication problem. (22/45) x (25/14) = 55/63 think of it the same as doing (1/2) * (1/3) = 1/6

OpenStudy (anonymous):

o ok thanx